# Xinyuan Lu1    Liangming Pan 2    Yubo Ma3    Preslav Nakov4    Min-Yen Kan1

### v2

@@ 3.4 Model Training
As no prior work adopts the tool-augmented LLM framework for table reasoning, there does not exist training data to train the modules Table Formatter, Tool Maker and Explanation Generator ($\mathcal{F}$, $\mathcal{M}$, and $\mathcal{E}$) in TART. Previous studies have demonstrated that smaller LLMs can learn from distilling the generated outputs of larger teacher LLMs that have better reasoning capabilities [38, 34, 17]. Following this, we use a teacher LLM $\mathcal{L}$ to first synthesize tool-integrated solution trajectories for a set of seed table-based reasoning tasks. These high-quality solution trajectories serve as the blueprint from which we automatically extract and rearrange their components to build training sets for $\mathcal{F}$, $\mathcal{M}$, and $\mathcal{E}$, the three modules used in TART.
@@
For all modules, we use GPT-4 as the teacher LLM $\mathcal{L}$ to generate training data. As shown in Table 4, we select five diverse table reasoning datasets: two from TQA and three from TFV, spanning general knowledge (Wikipedia) as well as domains such as finance, health, and scientific documents. These datasets provide a broad range of reasoning types. We few-shot prompt $\mathcal{L}$ to generate tool-integrated solutions for training instances for each dataset. In each solution, the model is prompted to clean the table, invent useful tools, and propose a reasoning plan along with explanations. We provide our prompt in Appendix F.
After generating the solutions, we evaluate the final answers against the ground truth, retaining only the instances with correct answers. Subsequently, we refine the solutions by removing overly specific tools through tool abstraction and tool deduplication. Tool abstraction filters out tools that appear only once, keeping those with broader applicability. Tool deduplication consolidates similar tools that perform the same function, but have different names or implementations (e.g., add and sum). As a result, we obtain 11,701, 9,916, and 9,916 training instances for the table formatter $\mathcal{F}$, tool maker $\mathcal{M}$, and explanation generator $\mathcal{E}$, respectively. We refer to this training dataset as ToolTab, with detailed statistics provided in Table 5.
@@
Instruction fine-tuning [24, 10] has emerged as a critical strategy that directs LLMs to adhere to specified instructions, facilitating their reasoning capability across a wide range of table-based tasks. Therefore, we use open-source LLMs with instruction tuning as the backbone models for the modules of TART, specifically Llama2-7B [32], Llama3-8B, CodeLlama-7B [30] and Deepseek-Coder-7B-Instruct-V1.5 [14]. We fine-tune all TART modules independently on their respective training datasets from ToolTab, using the standard next-token prediction objective.
@@
Table 1: Performance evaluation across backbone models using the TART framework, highlighting the best (bold) and second-best (underlined) results. The accuracy is calculated on testing sets, with overall average accuracy in the last column (Avg. Acc.). The red number indicates the average increase percentage over the CoT methods.
@@ 4 Experiments
Datasets and Baselines
To rigorously evaluate the performance of our proposed TART framework, we select two categories of benchmarks for table-based reasoning. (1) Table question answering (TQA) benchmarks: WikiTableQuestion (WTQ) [28] focuses on simple factoid questions. HiTab (HIT) [9], TabMWP (TMP) [21] and FinQA (FQA) [8] datasets focus on numerical reasoning reasoning. TAT-QA (TAT) [48] and HybridQA (HYQ) [7] require joint reasoning over the table and the text for financial reports and Wikipedia tables, respectively. 2) Table-based fact verification (TFV): We select TabFact (TAF) [6], SCITAB (SCT) [22], and PubHealthTab (PHT) [1] datasets, which focus on verifying facts based on tables from Wikipedia, scientific articles, and public health articles, respectively.
For baseline comparisons, we select well-known table-based open-source LLMs such as TableLlama [44] , as well as text-pretrained models (Llama2-7b, Llama3-8b) and code-pretrained models (CodeLlama-7b, DeepSeek-Coder-7b). We choose the 7b and the 8b versions to represent a balance between computational efficiency and the capacity for complex reasoning and generalization. For each model, we fine-tune with two settings: (1) DirectQA, where models generate answers directly from questions and tables, and (2) Chain-of-Thought (CoT) reasoning, which requires models to formulate a step-by-step reasoning process before concluding with an answer.
Implementation
@@ 4.1 Main Results
4. Using closed-source models (GPT-3.5-turbo and GPT-4) as the backbone models for TART achieves an average accuracy of 74.9, significantly outperforming the open-source counterparts, which average at 60.0 accuracy. Nonetheless, the highest-performing open-source model DeepSeek-7b reaches up to 90.0% of GPT-3.5-turbo’s performance and 77.6% of GPT-4, illustrating the competitiveness of open-source models in the creation and use of tools despite the model size gap.
@@ 4.2 Out-of-Domain Results
We hypothesize that the TART has enhanced generalization capabilities compared to CoT due to its ability to create and use general tools. To validate this, we further evaluate TART across four different out-of-domain (OOD) datasets: HiTab (HIT), WikiTableQuestion (WTQ), TAT-QA (TAT), and HybridQA (HYQ). The results are shown in Table 2.
The TART method demonstrates variable effectiveness, with notable improvements in certain contexts. For instance, it achieves an average accuracy increase of 29.3% on the WTQ dataset, indicating robust domain-transfer capabilities. The Deepseek-7b backbone model particularly excels, with 30.6% increase in accuracy. We hypothesize that this superiority stems from its pre-training on coding tasks, which equips it with the capability of effectively creating and using tools in novel domains, surpassing pure-text-based pretraining models such as Llama-2-7b. The analysis in Section 4.4 supports our hypothesis, suggesting that TART excels in developing generic table reasoning functions that generalize well across different domains.
Tab Formt TAF PHT SCT TMP FQA
Llama-2 71.8/78.5 75.8/66.4 64.0/57.0 93.6/92.0 73.4/37.7
Llama-3 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0
Codellama 67.6/78.0 81.2/66.1 64.6/53.9 94.1/91.5 76.1/35.7
DeepSeek 70.7/79.7 72.5/71.3 63.5/51.3 95.7/93.9 74.5/38.6
Tool Mkr TAF PHT SCT TMP FQA
Llama-2 70.2/81.8 65.8/60.2 53.9/61.5 95.7/91.1 61.7/31.0
Llama-3 75.5/75.4 71.1/69.8 63.5/52.2 97.9/92.4 62.2/38.5
Codellama 75.5/85.2 74.5/71.2 62.9/57.1 95.7/91.7 68.1/39.8
Deepseek 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0
Table 3: Results of TART with different backbone modules. The top section uses deepseek-code-7b as the Tool Maker, while the bottom section uses Llama-3-8b as the Table Formatter. The best performance is highlighted in bold and the second-best is underlined. Tab Formt stands for Table Formatter and Tool Mkr for Tool Maker.
4.3 Impact of Foundation Models
To explore the optimal module combinations within the TART framework, we explore various pairings of table formatter and toolmaker modules shown in Table 3. We find that using Llama-3-8B as the table formatter and DeepSeek-7B as the tool maker achieves the best average execution rate (76.8) and accuracy (68.6). This aligns with our expectations given that Llama-3-8B excels in processing long tables while DeepSeek-7B, with its pre-training on code, demonstrates superior capability in tool creation. Detailed results are shown in Table 6 in Appendix B.
Rank Tool Name
1 get_column_by_name
2 get_column_cell_value
3 get_row_index_by_value
4 extract_price
5 get_row_by_name
6 equal_to
7 get_column_by_index
8 subtract
9 divide
10 add
Figure 3: Analysis of tool usage in our TART framework. Shown are the distribution (left) and the categories of the top-20 tools across models (center), as well as the top-10 tools in TART Codellama-7b (right).
@@ 4.4 Analysis of Tool Creation
We then performed an in-depth analysis of how TART creates and utilizes tools shown in the Section 4.4.1 and Section 4.4.2.
@@
Figure 3 (left) illustrates the tool usage distribution across different backbone models in TART, highlighting a long-tail distribution. The most frequently used tools are primarily associated with table processing (e.g., get_column_by_name) and numerical reasoning (e.g., add), aligning with our observations in Section 4.1. Figure 3 (right) provides a detailed breakdown of tool categories for the top-30 tools, showing that table preprocessing and numerical reasoning tools are the most prevalent. This supports the consistency of tool utilization patterns within TART.
@@
Figure 4: Tool overlap distribution between InD and OOD datasets across different models.
Figure 5: Comparison of the number of repeat tools and correct tools across different models.
@@
Figure 5 further reveals that although Llama2-7b frequently reuses tools, it often applies them inappropriately. In contrast, CodeLlama-7b not only exhibits a high rate of tool reuse, but also demonstrates a greater accuracy in their appropriate application. Meanwhile, Llama3-8b, despite its lower rate of tool reuse, excels in the correct usage of tools, which contributes to its superior performance.
Figure 6: Case study of TART comparing it to CoT reasoning. Panel (a) illustrates a numerical calculation error in CoT where incorrect arithmetic leads to a wrong answer, and panel (b) demonstrates a table location error where CoT fails to retrieve the correct table values. Both errors can be reduced by TART through tool integration.
@@ 4.5 Case Study
To gain deeper insights into the advantages of TART over CoT reasoning, we conducted a case study, shown in Figure 6. The examples highlight the limitation of CoT in numerical reasoning and table preprocessing, such as incorrect calculation (Figure 6 a) and incorrect retrieval (Figure 6 b). Conversely, TART overcomes these challenges effectively via integrating specialized tools like subtract and get_column_by_index. Despite these strengths, TART still encounters issues related to data type mismatches and incorrect programming plans. A detailed analysis of error types in TART can be found in Appendix E.
5 Conclusion and Future Work
In this paper, we introduce an open-source framework to improve table-based reasoning through the Tool-Augmented Reasoning framework for Tables (TART). This framework solves the challenges of current LLMs’ limited ability to understand table structure and execute precise numerical calculations, and maintains explainability. TART consists of a table formatter for accurate data representation, a tool maker for creating specialized tools, and an explanation generator maintaining interpretable explanations. To train TART, we present the ToolTab dataset, a novel benchmark containing a diverse set of real-world tables and their tool-augmented solutions. Experiments across nine benchmarks show that integrating our TART method into different open-sourced LLMs enhances accuracy on table-based reasoning. Furthermore, in-depth analysis revealed that TART effectively learns and uses tools. Future work could extend TART to a multimodal framework by incorporating image-based question-answering and fact-verification to generate richer explanations. Additionally, generating explanations to satisfy the needs of different end users, such as laypeople and experts, could further improve the TART’s applicability and impact.
Limitation
@@ 5 Conclusion and Future Work
While our efforts have focused on expanding the range of our dataset to include a variety of tableQA datasets, some table-related datasets remain unrepresented in ToolTab. As a result, despite TART’s capacity to adapt to different OOD datasets and tasks, its performance might still different with the complexities and unique challenges of new table tasks and datasets that it has not yet encountered. Having initiated the development of an expansive, versatile tool-enhanced model for tables, we encourage continued research in this area to further advance the model’s ability to generalize across diverse table configurations.
Ethic Statement
@@
Our research uses datasets that are either publicly available or were collected with explicit consent. We adhere to strict data privacy and security protocols to protect the information and ensure it is used solely for the purposes of this research.
@@ Appendix A Dataset Composition for TART Training
In Table 4, we show the composition of the seed datasets used for training our TART model. These datasets vary in terms of the tasks, the domains, and the types of input and output data. For instance, TabMWP and FinQA focus on TableQA tasks within mathematics and finance domains respectively, requiring a combination of tables, text, and questions as inputs, with short answers as outputs. Meanwhile, PubHealthTab, TabFact, and SCITAB target table fact-checking tasks across health, general Wikipedia, and scientific article domains. These datasets similarly involve tables and claims as inputs but differ in the specifics of the domain-related claims, each producing a short label as an output. To construct the ToolTab, we obtain 11,701, 9,916, and 9,916 training instances for the table formatter $\mathcal{F}$, tool maker $\mathcal{M}$, and explanation generator $\mathcal{E}$, respectively. Detailed statistics are provided in Table 5.

### v3

@@ 3.4 Model Training
As no prior work adopts the tool-augmented LLM framework for table reasoning, there does not exist training data to train the modules Table Formatter, Tool Maker and Explanation Generator ($\mathcal{F}$, $\mathcal{M}$, and $\mathcal{E}$) in TART. Previous studies have demonstrated that smaller LLMs can learn from distilling the generated outputs of larger teacher LLMs that have better reasoning capabilities West et al. (2022); Wang et al. (2023); Kim et al. (2024). Following this, we use a teacher LLM $\mathcal{L}$ to first synthesize tool-integrated solution trajectories for a set of seed table-based reasoning tasks. These high-quality solution trajectories serve as the blueprint from which we automatically extract and rearrange their components to build training sets for $\mathcal{F}$, $\mathcal{M}$, and $\mathcal{E}$.
@@
For all modules, we use GPT-4 as the teacher LLM $\mathcal{L}$ to generate training data. As shown in Table 5, we select five diverse table reasoning datasets: two from TQA and three from TFV, spanning general knowledge (Wikipedia) as well as domains such as finance, health, and scientific documents. These datasets provide a broad range of reasoning types. We few-shot prompt $\mathcal{L}$ to generate tool-integrated solutions for training instances for each dataset. In each solution, the model is prompted to clean the table, invent tools, and propose a reasoning plan with explanations. We provide our prompt in Appendix F.
After generating the solutions, we evaluate the final answers against the ground truth, retaining only the instances with correct answers. Subsequently, we refine the solutions by removing overly specific tools through tool abstraction and tool deduplication. Tool abstraction filters out tools that appear only once, keeping those with broader applicability. Tool deduplication consolidates similar tools that perform the same function, but have different names or implementations (e.g., add and sum). As a result, we obtain 11,701, 9,916, and 9,916 training instances for the table formatter $\mathcal{F}$, tool maker $\mathcal{M}$, and explanation generator $\mathcal{E}$, We refer to this training dataset as ToolTab, with detailed statistics provided in Table 6.
@@
Instruction fine-tuning Mishra et al. (2022); Chung et al. (2024) has emerged as a critical strategy that directs LLMs to adhere to specified instructions, facilitating their reasoning capability across a wide range of table-based tasks. Therefore, we use open-source LLMs with instruction tuning as the backbone models for the modules of TART, specifically Llama-2-7b Touvron et al. (2023), Llama-3-8b, CodeLlama-7b Rozière et al. (2023) and Deepseek-Coder-7b-Instruct-V1.5 Guo et al. (2024). We fine-tune all TART modules independently on their respective training datasets from ToolTab, using the standard next-token prediction objective.
@@
Table 1: Performance evaluation across backbone models using the TART framework, highlighting the best (bold) and second-best (underlined) results. The accuracy is calculated on testing sets, with overall average accuracy in the last column (Avg. Acc.). The red numbers indicate the average increase percentage over the CoT methods.
@@ 4 Experiments
Datasets and Baselines.
To evaluate TART, we select two categories of benchmarks for table-based reasoning. (1) Table question answering (TQA): WikiTableQuestion (WTQ) Pasupat and Liang (2015) focuses on simple factoid questions. HiTab (HIT) Cheng et al. (2022), TabMWP (TMP) Lu et al. (2023b) and FinQA (FQA) Chen et al. (2021) datasets focus on numerical reasoning reasoning. TAT-QA (TAT) Zhu et al. (2021) and HybridQA (HYQ) Chen et al. (2020b) require joint reasoning over the table and the text for financial reports and Wikipedia tables, respectively. (2) Table-based fact verification (TFV): We select TabFact (TAF) Chen et al. (2020a), SCITAB (SCT) Lu et al. (2023c), and PubHealthTab (PHT) Akhtar et al. (2022) datasets, which focus on verifying facts based on tables from Wikipedia, scientific articles, and public health articles, respectively.
For baseline comparisons, we select well-known table-based open-source LLMs such as TableLlama Zhang et al. (2024a) , as well as text-pretrained models (Llama2-7b, Llama3-8b) and code-pretrained models (CodeLlama-7b, DeepSeek-Coder-7b). We choose the 7b and the 8b versions to represent a balance between computational efficiency and the capacity for complex reasoning and generalization. For each model, we fine-tune with two settings: (1) DirectQA, where models generate answers directly from questions and tables, and (2) Chain-of-Thought (CoT) reasoning, which requires models to formulate a step-by-step reasoning process before concluding with an answer.
Implementation.
@@ 4.1 Main Results
4. Using closed-source models (GPT-3.5-turbo and GPT-4) as the backbone models for TART achieves an average accuracy of 74.9, significantly outperforming the open-source counterparts, which average at 60.0 accuracy. Nonetheless, the highest-performing open-source model, DeepSeek-7b, reaches up to 90.0% of GPT-3.5-turbo’s performance and 77.6% of GPT-4, illustrating the competitiveness of open-source models in the creation and use of tools despite the apparent model size gap.
Model TAF PHT SCT TMP FQA Avg. Acc.
Llama2 69.2 55.0 53.4 88.8 19.2 57.1
- TabFT 67.7 39.1 50.4 67.8 17.2 48.4 (-15.2%)
[…]
- TabFT 69.2 58.9 46.3 65.4 13.8 50.7 (-16.9%)
[…]
- TabFT 66.6 64.9 36.6 69.1 20.2 51.5 (-13.2%)
[…]
- TabFT 58.0 58.8 45.0 71.3 16.7 50.0 (-19.9%)
Table 2: The ablation study of Table Formatter (- TabFT) in TART with different backbone models. The red numbers indicate the average accuracy (Avg. Acc.) drop percentage without Table Formatter for each model.
Figure 3: Ablation Study of TART’s Tool Maker and Explanation Generator. Panel (a) compares the pure program with TART’s tool-augmented program, highlighting the effectiveness of Tool Maker. Panel (b) compares TART’s tool-augmented program with TART’s final answer, demonstrating the usefulness of Explanation Generator.
4.2 Ablation Study
Ablation of the Table Formatter.
We conduct ablation experiments across all backbone models (Table 2). TART without the Table Formatter led to significant and uniform performance drops of over 10%. This clearly demonstrates the effectiveness of the Table Formatter module in improving reasoning capabilities by ensuring consistent table representations.
Ablation of the Tool Maker.
In this ablation, the framework directly generates programs instead of creating modular tools (Figure 3a). While this approach achieves functionality, it impacts both performance and explanability. These tools (e.g., get_column_by_name) can be used repeatedly comparing to the tools that are overly specific (e.g.,calculate_cookie_difference).
Ablation of the Explanation Generator.
The Explanation Generator module does not directly impact the final performance in terms of accuracy, but enhances the explanability of the outputs. For example, in Figure 3(b), TART’s final answer provides more human-readble explanations, comparing to the program.
Tab Formt TAF PHT SCT TMP FQA
Llama2 71.8/78.5 75.8/66.4 64.0/57.0 93.6/92.0 73.4/37.7
Llama3 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0
CodeLlama 67.6/78.0 81.2/66.1 64.6/53.9 94.1/91.5 76.1/35.7
DeepSeek 70.7/79.7 72.5/71.3 63.5/51.3 95.7/93.9 74.5/38.6
Tool Mkr TAF PHT SCT TMP FQA
Llama2 70.2/81.8 65.8/60.2 53.9/61.5 95.7/91.1 61.7/31.0
Llama3 75.5/75.4 71.1/69.8 63.5/52.2 97.9/92.4 62.2/38.5
CodeLlama 75.5/85.2 74.5/71.2 62.9/57.1 95.7/91.7 68.1/39.8
Deepseek 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0
Table 3: Results of TART with different backbone modules. The top half uses deepseek-code-7b as the Tool Maker, while the bottom half uses Llama3-8b as the Table Formatter. The best performance is highlighted in bold and the second-best is underlined. Tab Formt is Table Formatter and Tool Mkr is Tool Maker.
4.3 Impact of Foundation Models
To explore the optimal module combinations within the TART framework, we explore various pairings of table formatter and toolmaker modules shown in Table 3. We find that using Llama3-8b as the table formatter and DeepSeek-7b as the tool maker achieves the best average execution rate (76.8) and accuracy (68.6). This aligns with our expectations given that Llama3-8b excels in processing long tables while DeepSeek-7b, with its pre-training on code, demonstrates superior capability in tool creation (c.f. Table 7 in Appendix B).
5 Discussion
To further explore the usefulness of TART, we focus on the following research questions: 1) What is the performance of TART on out-of-domain (OOD) datasets? (Section 5.1); 2) How does TART create and utilize tools? (Section 5.2); and 3) What is the quality of the explanations generated by TART in actual user cases? (Section 5.3).
@@ 5.1 Out-of-Domain Results
Rank Tool Name
1 get_column_by_name
2 get_column_cell_value
3 get_row_index_by_value
4 extract_price
5 get_row_by_name
6 equal_to
7 get_column_by_index
8 subtract
9 divide
10 add
Figure 4: Analysis of tool usage in the TART framework. Th left and center panels show the distribution and the categories of the top 20 most-used tools across models. The right shows the top-10 tools in TART CodeLlama-7b.
@@
We hypothesize that the TART has enhanced generalization capabilities compared to CoT due to its ability to create and use general tools. To validate this, we further evaluate TART across four different OOD datasets: HiTab (HIT), WikiTableQuestion (WTQ), TAT-QA (TAT), and HybridQA (HYQ). The results are shown in Table 4.
The TART method demonstrates variable effectiveness, with notable improvements in certain contexts. For instance, it achieves an average accuracy increase of 29.3% on the WTQ dataset, indicating robust domain-transfer capabilities. The Deepseek-7b backbone model excels, with 30.6% increase in accuracy. We hypothesize that this superiority stems from its pre-training on coding tasks, which equips it with the capability of effectively creating and using tools in novel domains, surpassing pure-text-based pretraining models such as Llama2-7b. The analysis in Section 5.2 supports our hypothesis, suggesting that TART excels in developing generic table reasoning functions that generalize well across various domains.
@@ 5.2 Analysis of Tool Creation
We then performed an in-depth analysis of how TART creates and utilizes tools.
@@
Figure 4 (top-left) illustrates the tool usage distribution across different backbone models in TART, highlighting a long-tail distribution. The most frequently used tools are primarily associated with table processing (e.g., get_column_by_name) and numerical reasoning (e.g., add), aligning with our observations in Section 4.1. Figure 4 (middle) provides a detailed breakdown of tool categories for the top 30 tools, showing that table preprocessing and numerical reasoning tools are the most prevalent. This supports the consistency of tool utilization patterns within TART.
Figure 5: Case Study of TART comparing with CoT reasoning. Panel (a) illustrates a numerical calculation error in CoT where incorrect arithmetic leads to a wrong answer, and panel (b) demonstrates a table location error where CoT fails to retrieve the correct table values. Both errors can be reduced by TART through tool integration.
@@
Figure 6: Tool overlap distribution between in-domain and OOD datasets across different backbone models.
Figure 7: Comparison of the number of repeat tools and correct tools across different backbone models.
@@
Figure 7 further reveals that although Llama2-7b frequently reuses tools, it often applies them inappropriately. In contrast, CodeLlama-7b not only exhibits a high rate of tool reuse but also demonstrates a greater accuracy in their appropriate application. Meanwhile, Llama3-8b, despite its lower rate of tool reuse, excels in the correct usage of tools, contributing to its superior performance.
@@ 5.3 Case Study
To gain deeper insights into the advantages of TART over CoT reasoning, we conducted a case study, shown in Figure 5. The examples highlight the limitation of CoT in numerical reasoning and table preprocessing, such as incorrect calculation in Figure 5(a) and incorrect retrieval inFigure 5(b). Conversely, TART overcomes these challenges effectively via integrating specialized tools like subtract and get_column_by_index. Despite these strengths, TART still encounters issues related to data type mismatches and incorrect programming plans. A detailed analysis of error types in TART can be found in Appendix E.
6 Conclusion
We introduce an open-source framework to improve table-based reasoning through the Tool-Augmented Reasoning framework for Tables (TART). This framework solves the challenges of current LLMs’ limited ability to understand table structure and execute precise numerical calculations, and maintains explainability. TART consists of a table formatter for accurate data representation, a tool maker for creating specialized tools, and an explanation generator maintaining interpretable explanations. To train TART, we present the ToolTab dataset, a novel benchmark containing a diverse set of real-world tables and their tool-augmented solutions. Experiments across nine benchmarks show that integrating our TART method into different open-sourced LLMs enhances accuracy on table-based reasoning. Furthermore, in-depth analyses reveal that TART effectively learns and uses tools. Future work could extend TART to a multimodal framework by incorporating image-based question-answering and fact-verification to generate richer explanations. Additionally, generating explanations to satisfy the needs of different end users, such as laypersons and experts, could further improve the TART’s applicability and impact.
Limitations
@@ 6 Conclusion
While our efforts have focused on expanding the range of our dataset to include a variety of tableQA datasets, some table-related datasets remain unrepresented in ToolTab. As a result, despite TART’s capacity to adapt to different OOD datasets and tasks, its performance might still different with the complexities and unique challenges of new table tasks and datasets that it has not yet encountered. Having initiated the development of an expansive, versatile tool-enhanced model for tables, we encourage for continued research in this area to further advance the model’s ability to generalize across diverse table configurations.
Ethics Statement
@@
Our research utilizes datasets that are either publicly available or collected with explicit consent. We adhere to strict data privacy and security protocols to protect the information and ensure it is used solely for the purposes of this research.
@@ Appendix A Dataset Composition for TART Training
In Table 5, we show the composition of the seed datasets utilized for training our TART model. These datasets vary in terms of the tasks, the domains, and the types of input and output data. For instance, TabMWP and FinQA focus on TableQA tasks within mathematics and finance domains respectively, requiring a combination of tables, text, and questions as inputs, with short answers as outputs. Meanwhile, PubHealthTab, TabFact, and SCITAB target table fact-checking tasks across health, general Wikipedia, and scientific article domains. These datasets similarly involve tables and claims as inputs but differ in the specifics of the domain-related claims, each producing a short label as an output.